OpenStudy (anonymous):

If f(x)=√(6−5x), find a formula for f^−1(x).

sam (.sam.):

let y=√(6−5x) y^2=(6-5x) y^2-6=-5x x=(6-y^2)/5 f^-1(x)=(6-x^2)/5